OpenStudy (anonymous):

Each of the 75 children in a line was assigned one of the integers from 1 through 75 by counting off in order. Then, standing in the same order, the children counted off in the opposite direction, so that the child who was assigned the number 75 the first time was assigned the number 1 the second time. Which of the following is a pair of numbers assigned to the same child? A) 50 and 25 B) 49 and 24 C) 48 and 26 D) 47 and 29 E) 45 and 32

OpenStudy (ganpat):

well m not sure whether this is a standard formula or not.. but if u check some of the numbers it will come out to be 76.. eg.. for 1st student = 1 + 75 = 76 2nd student = 2 + 74 = 76.. and so on.. So can u calculate now ??

OpenStudy (ganpat):

1. 50 + 25 = 75 2. 49 + 24 = 75 3. 47 + 29 = 76 4. 45 + 32 = 77. so the answer is 3 rd option.. 47 + 29..

OpenStudy (anonymous):

is that the only way to solve it other than writing all the numbers out ofcourse

OpenStudy (ganpat):

i am not sure.. i need to think a little more.. but this give u answer and is one of the shortest way, dont u think so ??

OpenStudy (anonymous):

oh no i was just curious but ya this is a good way to do it thank you :D

OpenStudy (ganpat):

w/c..